# Cross-Lingual and Cross-Cultural Variation in Image Descriptions

### v1

@@ 1 Introduction
Figure 1: An image taken in an Indonesian-speaking area, corresponding captions in Indonesian (id) and Dutch (nl) translated to English, and saliency for the person.n.01 ($SAL_{p}$) and dancer.n.01 ($SAL_{d}$) synsets. Saliency is measured as the proportion of captions referring to the synset or any of its descendants (e.g., dancer.n.01 is a descendant of person.n.01).
Do speakers of different languages talk differently about what they see? To answer this question, we can compare their descriptions of the same images. In this setting, differences in semantic content must stem from linguistic rather than perceptual differences. [^2] It must be noted that Mitterer et al. (2009) provide evidence of environmental effects on visual perception; however, these effects are limited to low-level details, such as similar colour discrimination. Crucially, this takes into account certain aspects of culture (e.g., the saliency of an entity) while excluding environmental ones (e.g., the presence of an entity) (Liu et al., 2021; Hershcovich et al., 2022).
Previous work investigated which cultures draw more attention to the foreground or the context of a visual scene Miyamoto et al. (2006), and how often each culture uses negation or certain bigrams at the start of the descriptions van Miltenburg et al. (2017). However, these studies were conducted on a small scale, either with artificial images or focused on just a few related languages. To induce more solid conclusions, a large dataset of image descriptions in a variety of typologically and areally diverse languages is needed (Ponti et al., 2019).
A dataset with these properties has recently been released, named Crossmodal3600 (XM3600, Thapliyal et al., 2022). The XM3600 dataset is both relatively large in scale (3600 images) compared with previous controlled studies and more linguistically diverse (36 languages) compared with any previous image-description dataset. We propose to exploit these unique features to study cross-cultural differences in image descriptions.
Specifically, we focus on the entities that the annotators chose to mention in their descriptions. We develop automatic tools for accurately identifying phrases representing entities in image descriptions and map these phrases into WordNet Miller (1992) synsets, which function as a proxy for entity categories. We further annotate images with the categories they contain, independent of the descriptions. Finally, we study how entity mentions vary across different languages (see Figure 1 for an example).
Using this collected dataset of cross-lingual entity mentions as quantitative evidence, we are able to address the following questions:
@@
Our method can be used both as an exploration and a verification tool. For the former, we present new findings regarding cultural effects on object saliency (Sections 4.2, 4.3, and 4.4). Future work may propose small-scale, controlled studies to verify these findings.
@@
To summarise, our contributions are as follows. First, we develop an accurate methodology for identifying entities mentioned in image captions. Second, we present new insights into cross-lingual variation in entity saliency and confirm previous findings on a wider and more diverse set of languages. Lastly, we release our entity mention dataset to support future research on cross-lingual and cross-cultural variation of image descriptions.
@@ 3.1 Captions Translation
We first translate captions from all languages into English. Translation is a necessity if we are to consider a wide variety of languages since the tools required to automatically process the captions (e.g., WordNet and part-of-speech taggers) are only available in a few languages. Nevertheless, the translation process presents certain benefits, notably the elimination of metaphorical expressions containing entities in non-English languages, and thus a possible confounding factor. We use the Google Translation API [^3] https://cloud.google.com/translate following Thapliyal et al. (2022).
We find that translation quality was poor for the 5 low-resource languages [^4] Bengali (bn), Maori (mi), Cusco Quechua (quz), Swahili (sw), Telugu (te). of XM3600, as demonstrated by Benjamin (2019), where captions for these languages were marked as correct by native speakers 4 times less frequently. We therefore focus on the remaining 31 languages. [^5] Arabic (ar), Chinese-Simplified (zh), Croatian (hr), Czech (cs), Danish (da), Dutch (nl), English (en), Filipino (fil), Finnish (fi), French (fr), German (de), Greek (el), Hebrew (he), Hindi (hi), Hungarian (hu), Indonesian (id), Italian (it), Japanese (ja), Korean (ko), Norwegian (no), Persian (fa), Polish (pl), Portuguese (pt), Romanian (ro), Russian (ru), Spanish (es), Swedish (sv), Thai (th), Turkish (tr), Ukrainian (uk), Vietnamese (vi).
@@ 4.2 Geographic and Genetic Effects
We use lang2vec Littell et al. (2017) to source pre-computed geographical, genetic and featural distances between languages.

### v2

@@ 1 Introduction
Translated captions Saliency
person.n.01 dancer.n.01
id A dancer in a traditional Balinese costume…
Two dancers dressed in traditional Balinese… $\frac{2}{2}=1$ $\frac{2}{2}=1$
nl Asian traditional costume during a performance
Two people in costume for a performance
Traditional Balinese dancers in costumes $\frac{2}{3}$ $\frac{1}{3}$
Figure 1: A photo taken in an Indonesian-speaking area, corresponding captions in Indonesian (id) and Dutch (nl) translated to English, and saliency for the person.n.01 and dancer.n.01 synsets. Saliency is measured as the proportion of captions referring to the synset or any of its descendants (e.g., dancer.n.01 is a descendant of person.n.01).
Do speakers of different languages talk differently about what they see? This question is important for both machine learning and cognitive science. From a computational perspective, the answer can guide the creation of more diverse vision-and-language datasets (Ye et al., 2023). From a cognitive science perspective, it can help us study cultural effects on perception and language.
To investigate their cross-lingual and cross-cultural variation, we can compare the descriptions of the same images by different communities of speakers. In this setting, differences in semantic content must stem from linguistic rather than perceptual differences. [^2] It must be noted that Mitterer et al. (2009) provide evidence of environmental effects on visual perception; however, these effects are limited to low-level details, such as similar colour discrimination. Crucially, describing the same image takes into account cultural effects (e.g., the saliency of an entity) while excluding environmental ones (e.g., the presence of an entity) (Liu et al., 2021; Hershcovich et al., 2022).
Previous work investigated which cultures draw more attention to the foreground or the context of a visual scene Miyamoto et al. (2006), and how often each culture uses negation or certain bigrams at the start of the descriptions van Miltenburg et al. (2017). Nevertheless, these studies were conducted on a small scale, either with artificial images or focused on just a few related languages.
To induce more solid conclusions, a large dataset of image descriptions in a variety of typologically and areally diverse languages is needed (Ponti et al., 2019). We use Crossmodal3600 (XM3600, Thapliyal et al., 2022), which is both relatively large in scale (3600 images) compared with previous controlled studies and more linguistically diverse (36 languages) compared with any previous image–description dataset.
We develop automatic tools for accurately identifying phrases mentioning entities in image descriptions and map these phrases into WordNet Miller (1992) synsets, which function as a proxy for entity categories. We further annotate images with the categories they contain, independent of the descriptions. Finally, we study how entity mentions vary across different languages (see Figure 1 for an example).
Using this annotation as quantitative evidence, we are able to address the following questions:
@@
Our method can be used both as an exploration and a verification tool for cultural effects on object saliency. For the former, we present new findings (Sections 4.2, 4.3, and 4.4). Future work may propose small-scale, controlled studies to verify these findings.
@@
To summarise, our contributions are as follows. First, we present new insights into cross-lingual variation in entity saliency and confirm previous findings on a wider and more diverse set of languages. Second, we release our entity mention dataset to support future research on cross-lingual and cross-cultural variation of image descriptions.
@@ 3.1 Captions Translation
We first translate captions from all languages into English. Translation is a necessity if we are to consider a wide variety of languages since the tools required to automatically process the captions (e.g., WordNet and part-of-speech taggers) are only available in a few languages. Nevertheless, the translation process presents certain benefits, notably the elimination of metaphorical expressions containing entities in non-English languages, and thus a possible confounding factor.
Translation quality plays a critical role in our experiments, as incorrect translations can impact our conclusions. We therefore use the Google Translation API [^3] https://cloud.google.com/translate, which demonstrates robust performance in 31 out of the 36 languages in XM3600 according to several studies Jiao et al. (2023); Enis and Hopkins (2024). [^4] Arabic (ar), Chinese-Simplified (zh), Croatian (hr), Czech (cs), Danish (da), Dutch (nl), English (en), Filipino (fil), Finnish (fi), French (fr), German (de), Greek (el), Hebrew (he), Hindi (hi), Hungarian (hu), Indonesian (id), Italian (it), Japanese (ja), Korean (ko), Norwegian (no), Persian (fa), Polish (pl), Portuguese (pt), Romanian (ro), Russian (ru), Spanish (es), Swedish (sv), Thai (th), Turkish (tr), Ukrainian (uk), Vietnamese (vi). We exclude the remaining 5 low-resource languages [^5] Bengali (bn), Maori (mi), Cusco Quechua (quz), Swahili (sw), Telugu (te). from our experiments, as the Google Translation API is rated poorly by native speakers of these languages Benjamin (2019).
@@ 4.2 Geographic and Genetic Effects
We use lang2vec Littell et al. (2017) to source pre-computed geographical, genetic and featural distances between languages. Featural distance is the overlap between phonemic inventories.